# Temporal Scaling Law for Large Language Models

### v2

@@ Abstract
Recently, Large Language Models (LLMs) have been widely adopted in a wide range of tasks, leading to increasing attention towards the research on how scaling LLMs affects their performance. Existing works, termed Scaling Laws, have discovered that the final test loss of LLMs scales as power-laws with model size, computational budget, and dataset size. However, the temporal change of the test loss of an LLM throughout its pre-training process remains unexplored, though it is valuable in many aspects, such as selecting better hyperparameters directly on the target LLM. In this paper, we propose the novel concept of Temporal Scaling Law, studying how the test loss of an LLM evolves as the training steps scale up. In contrast to modeling the test loss as a whole in a coarse-grained manner, we break it down and dive into the fine-grained test loss of each token position, and further develop a dynamic hyperbolic-law. Afterwards, we derive the much more precise temporal scaling law by studying the temporal patterns of the parameters in the dynamic hyperbolic-law. Results on both in-distribution (ID) and out-of-distribution (OOD) validation datasets demonstrate that our temporal scaling law accurately predicts the test loss of LLMs across training steps. Our temporal scaling law has broad practical applications. First, it enables direct and efficient hyperparameter selection on the target LLM, such as data mixture proportions. Secondly, viewing the LLM pre-training dynamics from the token position granularity provides some insights to enhance the understanding of LLM pre-training.
@@
Yizhe Xiong1,2 Xiansheng Chen3 Xin Ye3 Hui Chen1,2 Zijia Lin3 Haoran Lian4Zhenpeng Su5Jianwei Niu4Guiguang Ding1,2 1Tsinghua University 2Beijing National Research Center for Information Science and Technology (BNRist) 3Kuaishou Technology 4Beihang University 5Institute of Information Engineering, Chinese Academy of Sciences, Beijing, China {xiongyizhe2001,yexin0807}@gmail.com chenxiansheng@pku.edu.cn huichen@mail.tsinghua.edu.cn {lianhaoran,niujianwei}@buaa.edu.cn linzijia07@tsinghua.org.cn dinggg@tsinghua.edu.cn suzhenpeng@iie.ac.cn
@@ 1 Introduction
Large Language Model (LLM) marks a paradigm shift in the scope of natural language processing, demonstrating unprecedented capabilities in accomplishing complicated tasks of natural language understanding and generation (Devlin et al., 2019; Radford et al., 2018). A cornerstone of this remarkable progress lies in the scalability of the Transformer architecture (Vaswani et al., 2017), which has facilitated the development of increasingly large models. Moreover, the accessibility of large-scale training is significantly enhanced by the abundance of data gathered from the Internet, which enables the construction of giant training datasets to improve model performance (Radford et al., 2019). Due to the scalability of both the model size and training data scale, LLMs with billions of parameters Brown et al. (2020) and even trillions of parameters Fedus et al. (2022) are widely proporsed and applied in various tasks.
@@
The scalability of LLMs has been extensively studied in terms of variables like model size, computational budget, and dataset size (Kaplan et al., 2020). Prior works in this domain, which are commonly referred to as “scaling laws”(Kaplan et al., 2020; Henighan et al., 2020), have proposed empirical principles to characterize the power-law (i.e., exponential patterns) among those variables. Those scaling laws demonstrate that the final test loss of an LLM improves exponentially with scalable factors, i.e., model size, computational budget, and dataset size. Recently, similar explorations have expanded to multi-modal training (Cherti et al., 2023; Aghajanyan et al., 2023) and transfer learning (Hernandez et al., 2021), enabling researchers to predict the performance of LLMs in different scenarios.
@@
Prior works of scaling laws typically predict the final test loss of fully-trained LLMs with a given computational budget after completing their pre-training with mostly fixed hyperparameters (e.g., data mixture proportions, weight decay, etc.) (Kaplan et al., 2020; Henighan et al., 2020; Cherti et al., 2023; Aghajanyan et al., 2023). However, variations in training hyperparameters also significantly influence the final test loss in a complicated way (Yang et al., 2022; Xie et al., 2023a). That underscores the need for a more fine-grained scaling law, which can predict the temporal evolution of the test loss as training steps scale up under different training hyperparameters on a fixed model size and data scale. Such a temporal scaling law is complementary to prior works, because it further enables the identification of better training hyperparameters, such as data configurations and learning rates, after the model size and data scale are determined based on previous research. Considering that many current LLMs are trained with hyperparameters selected on a much smaller one due to computational costs, our work addresses the recognized issue that hyperparameters like data mixture proportions (Xie et al., 2023b) and regularization hyperparameters (Yang et al., 2022) are not always transferrable from a small language model to a much larger one, via proposing a two-stage pipeline to directly select hyperparameters on the target LLM. Furthermore, such a temporal scaling law provides a theoretical basis for studying some training dynamics of LLMs. Despite the broad applications, prior works have not explored practical scaling laws from the temporal perspective.
@@
In this paper, we propose the novel concept of Temporal Scaling Law for Large Language Models. Specifically, we first try to model the test loss as a whole in a coarse-grained manner but find it to be not precise enough. Then we further break the test loss into the losses for tokens in different positions. And by careful investigation, we discovered that the losses on different token positions conform to a dynamic hyperbolic-law across different training steps. We further examine the evolution of the curve parameters for the dynamic hyperbolic-law and identify their temporal patterns. We term this phenomenon as the temporal scaling law, i.e., how the test loss of an LLM evolve as the training steps scale up. Our temporal scaling law accurately predicts the subsequent test losses using the data from an early training period. Prediction results on both the in-distribution (ID) and the out-of-distribution (OOD) validation datasets show that our methodology significantly improves over baseline approaches.
@@
Our temporal scaling law has broad practical applications for LLM pre-training. In this paper, we provide two use cases as examples. First, our temporal scaling law presents a novel and practical approach to selecting the hyperparameters directly on the target to-be-pre-trained LLM. Taking the data mixture proportion as an example, current works generally tune data proportions on a small-scale model and directly apply the tuned weights to pre-training the much larger target LLM (Xie et al., 2023a). However, optimal hyperparameters for a small-scale model probably are not the optimal ones for the much larger target LLM (Yang et al., 2022), though they can be effective candidates. With the accurate loss prediction from our proposed temporal scaling law, we can leverage a two-stage pipeline to better select the target LLM’s hyperparameters, i.e., firstly using a small model to narrow down the range of candidate hyperparameter values and then directly selecting the best one which can achieve the best predicted test loss after training the target LLM with it for a small-scale of training data. Second, our temporal scaling law reveals some learning dynamics of LLMs at the token granularity. Specifically, we discovered that the loss decrease rate for tokens on different positions remains uniform after an early training period. Through experiments on various position-based weighting strategies, we verify the effectiveness of the default pre-training practice, in which no weighting strategies on different token positions are applied, though they are imbalanced in terms of learning difficulty.
@@
  • First, we propose the Temporal Scaling Law for LLMs, modeling the evolution of test loss during the LLM pre-training process.
  • Second, based on our proposed temporal scaling law, we provide a method for precisely predicting the subsequent test losses after deriving the parameters of the temporal scaling law from an early training period.
  • Lastly, we present two scenarios to illustrate the broad practical applications of our temporal scaling law. For hyperparameter tuning, our temporal scaling law enables the direct selection of better hyperparameter values based on predicted LLM performance on the target LLM. For learning dynamics, our temporal scaling law provides theoretical and experimental verifications for the default pre-training practice that puts no weights on each token loss at different positions.
2 Related Works
2.1 Large Language Models
Language models are statistical models designed to model the probabilistic correlation in natural language sequences (Touvron et al., 2023a). The introduction of the Transformer architecture (Vaswani et al., 2017) led to the development of large language models. GPT-3 (Brown et al., 2020) marks the beginning of the LLM era, as decoder-based generative models are capable of completing various tasks via in-context learning. Further advancements in LLMs include PaLM (Chowdhery et al., 2023), Pythia (Biderman et al., 2023), LLaMA (Touvron et al., 2023a), etc. Currently, GPT-4 (OpenAI, 2023) has pushed the boundaries of LLMs further in terms of scale and capability.
2.2 Scaling Laws for Language Models
The concept of scaling laws for language models was proposed by (Kaplan et al., 2020). Their study revealed that the test loss for generative transformer models scales as a power-law with model size, dataset size, and the amount of computation used for training. Building upon that foundational study, further research has expanded the concept of scaling laws to diverse problem settings (Hernandez et al., 2021) and model architectures (Cherti et al., 2023; Aghajanyan et al., 2023). For instance,  (Hernandez et al., 2021) has investigated scaling laws for transfer learning. (Cherti et al., 2023) and (Aghajanyan et al., 2023) observed scaling laws for multi-modal model pre-training.
Despite previous advancements, a critical point that remains underexplored is the temporal trajectory of LLM performance throughout training. According to  (Kaplan et al., 2020), when LLMs are pre-trained with infinite data and training steps, the test loss follows the power-law. However, this assumption cannot be fulfilled in real-world, thus the power-law may not be accurate to portray the temporal pattern of test loss during actual pre-training. By studying the loss behavior on different token positions, we introduce temporal scaling law, allowing for more precise tracking and prediction of LLM test loss during pre-training.
@@ 3.1 Preliminaries
Pre-training Language Models. We mainly discuss the generation process of decoder-based generative language models. A generative language model is trained via next-token prediction. During the pre-training process, the text corpus is commonly segmented into token sequences with an identical length. The token sequences are then fed to the model for calculating the prediction loss on each token position $i$. Formally, given a training text sequence $T_{n}$ consisting of $n$ tokens, i.e., $T_{n}=[t_{1},\cdots,t_{i-1},t_{i},\cdots,t_{n}]$, when predicting the token $t_{i}$, the language model takes the previous tokens $T_{i-1}=[t_{1},\cdots,t_{i-1}]$ as input, and generates the probability distribution $\mathbf{p}_{i}$ for the token $t_{i}$. The loss $\mathcal{L}_{i}$ w.r.t. the token $t_{i}$ is commonly calculated via the cross-entropy loss function. In decoder-based transformers (Vaswani et al., 2017), a look-ahead mask is applied in the multi-head attention (MHA) module to ensure that each token can only attend to previous tokens, preserving the causal order of the generated sequence. Then, all tokens’ $\mathbf{p}_{i}$ are calculated in a single forward pass in parallel. The loss $\mathcal{L}_{i}$ on each token’s $\mathbf{p}_{i}$ is then averaged as the loss for the sequence $T_{n}$:
@@
3.2 Experiment Setup
We use experiment results to derive the temporal scaling law, and conduct predictions after deriving parameters of the temporal scaling law from an early training period. See Appendix for details.
Train Dataset. We train LLMs on the Pile (Gao et al., 2021), a large-scale, publicly available monolingual corpus in English for language model pre-training, consisting of 22 text domains. For all experiments, we tokenize it using the LLaMA tokenizer (Touvron et al., 2023a) with a 32k vocabulary. We apply the domain weights in (Su et al., 2023) for LLM pre-training.
Validation Dataset. We apply two validation datasets for test loss calculation. For the in-distribution (ID) dataset, we randomly sample 800 sequences (1024 consecutive tokens for each) for each text domain from the validation set in the Pile. For the out-of-distribution (OOD) dataset, we simply adopt the validation split of the RealNews-Like domain in the C4 dataset (Raffel et al., 2020), as it is claimed to be “distinct from the Pile” (Gao et al., 2021). We refer to both validation datasets as ID-Val and OOD-Val, respectively. When calculating test loss, we forward all sequences in the validation dataset and average the results. It is worth noting that calculating test loss is equivalent to calculating the perplexity (PPL) for the validation dataset since PPL can be directly acquired via test loss:
@@ 3.2 Experiment Setup
468M -2.8746 0.9988 -2.7967 0.9983
1.2B -2.3998 0.9995 -2.3780 0.9993
Table 1: $R^{2}$ results for fitting the test loss with the power-law and our temporal scaling law. Note that the fit is caculated on complete test loss data. The power-law fails in portraying the test loss evolution.
Models. We train two generative language models with 468M and 1.2B parameters, respectively. The architectures of models are identical to the 410M and 1.0B parameter models in (Biderman et al., 2023). The differences between model parameters can be ascribed to the different vocabulary sizes and different activation functions (i.e., SwiGLU v.s. GeLU). See Appendix for details.
@@
Metric for Fitting Model Evaluation. We propose the temporal scaling law to fit the test loss during pre-training. Following the common practical choice (Cheng et al., 2014), we choose the coefficient of determination ($R^{2}$) to evaluate the quality of the fit. $R^{2}$ demonstrates the proportion of the variability in the ground-truth that the proposed fit could explain. Formally, we have:
@@
where $y_{i}$ is the ground-truth test loss w.r.t. the $i$-th step, and $\overline{y}$ is the average of $y_{i}$. $f_{i}$ is the fit test loss w.r.t. the $i$-th step. $R^{2}\in(-\infty,1]$ and a perfect fit has $R^{2}=1$. A greater $R^{2}$ indicates a better fit.
@@ 3.3 Temporal Scaling Law
The original scaling law (Kaplan et al., 2020) has validated that the temporal loss evolution of LLMs during pre-training follows a power-law under the assumption that training tokens are infinite and the LLM is trained with infinite training steps. However, in real-world applications, the maximum step of the training schedule is limited. To validate the effectiveness of the power-law in a limited training schedule, we pre-train 468M and 1.2B models following Sec. 3.2, and fit the power-law function ($\mathcal{L}=(p_{1}x)^{p_{2}}+p_{3}$, $\{p_{i}\}$ are fitting parameters) to the test loss evolution curve (i.e., treating test loss as a whole to fit in this case) using non-linear least squares. As shown in Tab. 1, applying the power-law results in negative $R^{2}$ values, indicating that the power-law fails to portray the test loss evolution across model scales and validation datasets. Such results motivate us to find a better method to understand the test loss evolution under real-world limitations where training steps are limited during LLM pre-training.
The Dynamic Hyperbolic-Law of Loss on Different Token Positions. To fit and predict the test loss more precisely, we propose to break it down and investigate the temporal scaling law from a finer granularity and look into the inherent patterns of the loss on every token position. Language models are statistical models trained to model the probabilistic process of next-token prediction. Formally, for a sequence of $n$ tokens $T_{n}=(t_{1},t_{2},\cdots,t_{n})$, the probability of the following token being $t_{n+1}$ given the preceding tokens is denoted as $P(t_{n+1}|t_{1},t_{2},\cdots,t_{n})$. Intuitively and statistically, in a consecutive sequence:
@@
since $t_{n+1}$ has a longer context than $t_{n}$. To illustrate this pattern, we plot the results of test loss for tokens in different positions, as validated on our ID-Val, using both our 468M and 1.2B models after completing pre-training. As shown in  Fig. 1, both models, although varying in scales, follow the same pattern that test loss for tokens with shorter context is commonly higher than that for tokens with longer context. For both models, the loss values for tokens near the end of the sequence seem to converge at a certain value. We find that this trend can be fit with a hyperbolic relation:
@@
where $\mathcal{L}_{i}$ is the loss on token position $i(1\leq i\leq n)$, and $a_{0}$, $a_{1}$, and $a_{2}$ are fitting parameters solved by non-linear least squares. The fit for data from our 468M and 1.2B model are presented in Fig. 1. By applying this fit to test loss results of all checkpoints across the training phase, we find that over 99% of them can be fit with such hyperbolic relation with $R^{2}>0.95$, well demonstrating its generality. We term this phenomenon as the dynamic hyperbolic-law. The conclusion still holds when evaluated on the OOD-Val. We leave those results in the Appendix.
Figure 1: Graph of actual loss and fitting curve across different token positions for the 468M and 1.2B models on the ID-Val. The loss on different token positions follows a dynamic hyperbolic-law.
Figure 2: The fitting curve for $a_{0}$, $a_{1}$, and $a_{2}$ on ID-Val throughout the pre-training stage of the 468M model.
From the Dynamic Hyperbolic-law to the Overall Temporal Pattern. In the dynamic hyperbolic-law, $a_{2}$ is the converging factor denoting the converged value of loss on each token as context length increases. $a_{0}$ denotes the loss gap between the first token and the last token, and $a_{1}$ is the scaling factor along sequence length. To analyze the temporal pattern of the overall test loss, we collect the tuple $(a_{0},a_{1},a_{2})$ for all evaluated steps during pre-training phase: $[(a_{0}^{N_{1}},a_{1}^{N_{1}},a_{2}^{N_{1}}),\cdots,(a_{0}^{N_{tot}},a_{1}^{N_{tot}},a_{2}^{N_{tot}})]$, where $N_{i}$ is the number of trained tokens at step $i$, and $N_{tot}$ is the total number of tokens used for pre-training. We plot the original $(a_{0},a_{1},a_{2})$ data on ID-Val w.r.t. the number of tokens trained (denoted as $N$) for the 468M model in Fig. 2. We fit the temporal evolution of $a_{0}$, $a_{1}$, and $a_{2}$ factors to depict the overall temporal test loss pattern.
Empirically, we use the following functions to fit for $a_{0}$ and $a_{1}$:
@@
where the $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ fitting parameters are solved by non-linear least squares. Observed from the original data in Fig. 2, we find that the value of $a_{0}$ and $a_{1}$ generally converges after a period of training, but suffer from fluctuations after converging due to the uncertainty of the first prediction position in a sequence. To mitigate the fluctuations, we define a separation point ${N}_{sep}$:
@@
We manually stabilize the values of $a_{0}$ and $a_{1}$ after ${N}_{sep}$:

### v3

@@ Abstract
Recently, Large Language Models (LLMs) have been widely adopted in a wide range of tasks, leading to increasing attention towards the research on how scaling LLMs affects their performance. Existing works, termed Scaling Laws, have discovered that the final test loss of LLMs scales as power-laws with model size, computational budget, and dataset size. However, the temporal change of the test loss of an LLM throughout its pretraining process remains unexplored, though it is valuable in many aspects, such as selecting better hyperparameters directly on the target LLM. In this paper, we propose the novel concept of Temporal Scaling Law, studying how the test loss of an LLM evolves as the training steps scale up. In contrast to modeling the test loss as a whole in a coarse-grained manner, we break it down and dive into the fine-grained test loss of each token position, and further develop a dynamic hyperbolic-law. Afterwards, we derive the much more precise temporal scaling law by studying the temporal patterns of the parameters in the dynamic hyperbolic-law. Results on both in-distribution (ID) and out-of-distribution (OOD) validation datasets demonstrate that our temporal scaling law accurately predicts the test loss of LLMs across training steps. Our temporal scaling law has broad practical applications. First, it enables direct and efficient hyperparameter selection on the target LLM, such as data mixture proportions. Secondly, viewing the LLM pretraining dynamics from the token position granularity provides some insights to enhance the understanding of LLM pretraining.
@@
Yizhe Xiong1,2 Xiansheng Chen3 Xin Ye3 Hui Chen1,2 Zijia Lin3 Haoran Lian4 Zhenpeng Su5Wei Huang6Jianwei Niu4Jungong Han1Guiguang Ding1,2 1Tsinghua University 2Beijing National Research Center for Information Science and Technology (BNRist) 3Kuaishou Technology 4Beihang University 5Institute of Information Engineering, Chinese Academy of Sciences, Beijing, China 6School of Computer Science, Beijing University of Posts and Telecommunications xiongyizhe2001@gmail.com
@@ 1 Introduction
Large Language Model (LLM) marks a paradigm shift in the scope of natural language processing, demonstrating unprecedented capabilities in accomplishing complicated tasks of natural language understanding and generation Devlin et al. (2019); Radford et al. (2018). A cornerstone of this remarkable progress lies in the scalability of the Transformer architecture Vaswani et al. (2017), which has facilitated the development of increasingly large models. Moreover, the accessibility of large-scale training is significantly enhanced by the abundance of data gathered from the Internet, which enables the construction of giant training datasets to improve model performance Radford et al. (2019). Due to the scalability of both the model size and training data scale, LLMs with billions of parameters Brown et al. (2020) and even trillions of parameters Fedus et al. (2022) are widely proposed and applied in various tasks.
@@
The scalability of LLMs has been extensively studied in terms of variables like model size, computational budget, and dataset size Kaplan et al. (2020). Prior works in this domain, which are commonly referred to as “scaling laws”Kaplan et al. (2020); Henighan et al. (2020), have proposed empirical principles to characterize the power-law (i.e., exponential patterns) among those variables. Those scaling laws demonstrate that the final test loss [^1] Following Kaplan et al. (2020), ”test loss” refers to the loss on unseen data, which can include both the test and validation sets. of an LLM improves exponentially with scalable factors, i.e., model size, computational budget, and dataset size. Recently, similar explorations have expanded to multi-modal training Cherti et al. (2023); Aghajanyan et al. (2023) and transfer learning Hernandez et al. (2021), enabling researchers to predict the performance of LLMs in different scenarios.
Temporal Scaling Law Prior Scaling Laws
Predict objective Evolution of test loss during pretraining Final test loss after pretraining
Fitting & Predicting manners Fit with light early pretrain on target LLM, Fit with fully pretrained smaller models,
predict on the target LLM predict on the target LLM
Granularity Loss on each token position Loss as a whole
Application Select hyperparameters directly on target LLMs, Decide a target LLM size &
Provide insight w.r.t. pretraining dynamics, etc. data scale after given a budget
Table 1: Comparison between our temporal scaling law and prior scaling laws.
@@
Prior works of scaling laws typically predict the final test loss of fully-trained LLMs with a given computational budget after completing pretraining with mostly fixed hyperparameters (like data mixture proportions, weight decay, etc.) Kaplan et al. (2020); Henighan et al. (2020); Cherti et al. (2023); Aghajanyan et al. (2023). However, variations in training hyperparameters also significantly influence the final test loss in a complicated way Yang et al. (2022); Xie et al. (2023). That underscores the need for a more fine-grained scaling law, which is to predict the temporal evolution of the test loss as training steps scale up under different training hyperparameters on a fixed model size. Such a temporal scaling law is supposed to be complementary to prior works, because it further enables the direct identification of better training hyperparameters on the target LLM after the model size and data scale are determined based on previous research. Furthermore, such a temporal scaling law can also provide a theoretical basis for studying more training dynamics of LLMs. Despite broad applications, prior works have not explored practical scaling laws from the temporal perspective.
@@
In this paper, we propose the novel concept of Temporal Scaling Law for LLMs. Specifically, we first try to model the test loss as a whole in a coarse-grained manner but find it to be not precise enough. Then we further break the test loss into the losses for tokens in different positions. By carefully investigating multiple functions, we discovered that utilizing dynamic hyperbolic-law accurately portrays the pattern of losses on different token positions across different training steps. We further examine the evolution of the curve parameters for the dynamic hyperbolic-law and identify their temporal patterns. We term this phenomenon as the temporal scaling law, i.e., how the test loss of an LLM evolves as the training steps scale up. Our temporal scaling law accurately predicts the subsequent test losses using the data from an early training period. Prediction results on both the in-distribution (ID) and the out-of-distribution (OOD) validation datasets show that our methodology significantly improves over baseline approaches.
Our temporal scaling law has broad practical applications for LLM pretraining. In this paper, we provide two use cases as examples:
@@
(a) Our temporal scaling law presents a novel and practical approach to selecting the hyperparameters directly on the target to-be-pretrained LLM. Taking the data mixture proportion as an example, current works generally tune data proportions on a small-scale model and directly apply the tuned weights to pretraining the much larger target LLM Xie et al. (2023). However, optimal hyperparameters for a small-scale model probably are not the optimal ones for the much larger target LLM Yang et al. (2022). With the accurate loss prediction from our proposed temporal scaling law, we can select the target LLM’s hyperparameters directly by choosing the best one which can achieve the lowest-predicted test loss after training the target LLM with a small-scale of training data.
(b) Our temporal scaling law reveals some learning dynamics of LLMs at the token granularity. Specifically, we theoretically and experimentally discovered that the loss decrease rate for tokens on different positions remains uniform after an early training period. Through experiments on various position-based weighting strategies, we verify the effectiveness of the default pretraining practice, in which no weighting strategies on different token positions are applied, though they are imbalanced in terms of learning difficulty.
Differences from Existing Scaling Laws. We summarize the differences between our temporal scaling law and prior scaling laws Kaplan et al. (2020); Henighan et al. (2020); Cherti et al. (2023); Aghajanyan et al. (2023) in Tab. 1. Our temporal scaling law primarily focuses on the evolution of test loss during the pretraining process, while prior scaling laws model the relations of the final test loss and the computational budget (generally decided by model size and training data scale). Our temporal scaling law enables a direct selection of training hyperparameters on a target LLM after the model size and training data scale are decided by prior scaling laws. We underscore that our temporal scaling law is a generic pattern that decoder-based LLMs follow during the pretraining process.
@@
  • We propose the Temporal Scaling Law, modeling the evolution of test loss during LLM pretraining.
  • Based on our proposed temporal scaling law, we provide a method for precisely predicting the subsequent test losses after deriving the parameters of the temporal scaling law from an early training period.
  • We present two scenarios to illustrate the broad practical applications of our temporal scaling law. For hyperparameter tuning, our temporal scaling law enables the direct selection of better hyperparameter values based on predicted LLM performance on the target LLM. For learning dynamics, our temporal scaling law provides theoretical and experimental verifications for the default pretraining practice that puts no weight on each token loss at different positions.
@@ 2.1 Preliminaries
Pretraining Language Models. We mainly discuss the generation process of decoder-based generative language models. A generative language model is trained via next-token prediction. During pretraining, the text corpus is commonly segmented into token sequences. The token sequences are then fed to the model for calculating the prediction loss on each token position $i$. Formally, given a training text sequence $T_{n}$ consisting of $n$ tokens, i.e., $T_{n}=[t_{1},\cdots,t_{i-1},t_{i},\cdots,t_{n}]$, when predicting the token $t_{i}$, the language model takes the previous tokens $T_{i-1}=[t_{1},\cdots,t_{i-1}]$ as input, and generates the probability distribution $\mathbf{p}_{i}$ for the token $t_{i}$. The loss $\mathcal{L}_{i}$ w.r.t. the token $t_{i}$ is commonly calculated via the cross-entropy loss function. In decoder-based transformers Vaswani et al. (2017), a look-ahead mask is applied in the multi-head attention (MHA) module to ensure that each token can only attend to previous tokens, preserving the causal order of the generated sequence. Then, all tokens’ $\mathbf{p}_{i}$ can be calculated in a single forward pass in parallel during pretraining. The loss $\mathcal{L}_{i}$ on each token’s $\mathbf{p}_{i}$ is then averaged as the loss for sequence $T_{n}$:
@@
2.2 Experiment Setup for Deriving Temporal Scaling Law
We use experiment results to derive the temporal scaling law and conduct predictions after deriving parameters of the temporal scaling law from an early training period. Please refer to appendix B for more experiment details.
Train Dataset. We train LLMs on the Pile Gao et al. (2021), consisting of 22 English text domains. For all experiments, we tokenize it using the LLaMA tokenizer Touvron et al. (2023a) with a 32k vocabulary. We apply the domain weights in Su et al. (2023) for LLM pretraining.
Validation Dataset. We apply two validation datasets for test loss calculation. For the in-distribution (ID) dataset, we randomly sample 800 sequences (1024 consecutive tokens for each) for each text domain from the validation set in the Pile. For the out-of-distribution (OOD) dataset, we simply adopt the validation split of the RealNews-Like domain in the C4 dataset Raffel et al. (2020), as it is claimed to be “distinct from the Pile” Gao et al. (2021). We refer to both validation datasets as ID-Val and OOD-Val, respectively. When calculating test loss, we forward all sequences in the validation dataset and average the results. It is worth noting that calculating test loss is equivalent to calculating the perplexity (PPL) for the validation dataset since PPL can be directly acquired via test loss:
@@ 2.2 Experiment Setup for Deriving Temporal Scaling Law
9.8M 0.6319 0.9858 0.7052 0.9795
58M 0.7606 0.9961 0.7684 0.9954
Table 2: $R^{2}$ results for fitting the test loss with the power-law (treat test loss as a whole) and our temporal scaling law (token-level granularity). Bold represents the best result.
Models. We train two LLaMA-structure Touvron et al. (2023a) generative language models with 9.8M and 58M parameters to illustrate our temporal scaling law. The architectures of both are identical to the 14M and 70M parameter models in Biderman et al. (2023). The differences between model parameters can be ascribed to the different vocabulary sizes and different activation functions (i.e., SwiGLU v.s. GeLU). When utilizing the temporal scaling law for predictions and further applications, we adopt the architectures of the 410M and 1.0B parameter models in Biderman et al. (2023), which leads to 468M and 1.2B models, respectively. To further scale up and generalize the validation of temporal scaling law, we also pre-train: (1) a 6.7B model with the same architecture of LLaMA-7B, (2) a 468M GPT-NeoX model, and (3) an 2B MoE model. See appendix F for results on (2) and (3).
@@
Metric for Fitting Model Evaluation. We propose the temporal scaling law to fit the test loss during pretraining. Following the common practical choice Cheng et al. (2014), we choose the coefficient of determination ($R^{2}$) to evaluate the quality of the fit. $R^{2}$ demonstrates the proportion of the variability in the ground truth that the proposed fit could explain. Formally, we have:
@@
where $y_{i}$ is the ground-truth test loss w.r.t. the $i$-th step, and $\overline{y}$ is the average of $y_{i}$. $f_{i}$ is the fit test loss w.r.t. the $i$-th step. $R^{2}\in(-\infty,1]$ and a perfect fit has $R^{2}=1$. A larger $R^{2}$ indicates a better fit.
@@ 2.3 Temporal Scaling Law
The original scaling law Kaplan et al. (2020) has validated that the final test loss for training with different model sizes and data sizes follows a power-law. To validate the effectiveness of the power-law for the temporal evolution of test loss, we pretrain 9.8M and 70M models following Sec. 2.2 and fit the power-law function ($\mathcal{L}=(p_{1}x)^{p_{2}}+p_{3}$, $\{p_{i}\}$ are fitting parameters) to the test loss evolution curve (i.e., treating test loss as a whole to fit in this case) using non-linear least squares. As shown in Tab. 2, though somehow effective, applying the power-law is not precise enough to portray the temporal evolution for the test loss (follow-up experiments in Fig. 3 also demonstrate that directly using the power-law in predicting the future test loss results in larger errors). Such results motivate us to find a better method to understand the test loss evolution during LLM pretraining.
The Dynamic Hyperbolic-Law of Loss on Different Token Positions. To fit and predict the test loss more precisely, we propose to break it down and investigate the temporal scaling law from a finer granularity and look into the inherent patterns of the loss on every token position. Language models are statistical models trained to model the probabilistic process of next-token prediction. Formally, for a partial sequence of $n$ tokens $T_{n}=(t_{1},t_{2},\cdots,t_{n})$, the probability of the following token being $t_{n+1}$ given the preceding tokens is denoted as $P(t_{n+1}|T_{n})$. Intuitively and statistically, in a consecutive sequence:
@@
since $t_{n+1}$ has a longer context than $t_{n}$. To illustrate this pattern, we plot the results of test loss for tokens in different positions, as validated on our ID-Val, using both our 9.8M and 58M models after training for 100B, 200B, 300B, and 400B tokens. We show the results for 400B tokens and leave the rest in the Appendix. As shown in  Fig. 1, both models, although varying in scales, follow the same pattern that test loss for tokens with shorter context is commonly higher than that for tokens with longer context. For both models, the loss values for tokens near the end of the sequence seem to converge at a certain value. We find that this trend can be well fit with a hyperbolic relation:
@@
where $\mathcal{L}_{i}$ is the loss on token position $i(1\leq i\leq n)$, and $a_{0}$, $a_{1}$, and $a_{2}$ are fitting parameters solved by non-linear least squares. The fit for data from our 9.8M and 58M model are presented in Fig. 1. By applying this fit to test loss results of all checkpoints across the training phase, we find that over 99% of them can be fit with such hyperbolic relation with $R^{2}>0.95$, well demonstrating its generality. We term this phenomenon as the dynamic hyperbolic-law. See appendix C for more insights.
We also tested Kaplan’s power-law Kaplan et al. (2020) and other functions that conform to Eq. 4 to fit the curve, but the results are suboptimal. For example, using the power-law yields $R^{2}<0.75$ for over 95% checkpoints. The conclusion still holds on the OOD-Val and on larger-scale models. See detailed results in appendix E.
Figure 1: Actual loss and fitting curve across different token positions for the 9.8M and 58M models on ID-Val. The loss follows a dynamic hyperbolic-law.
Figure 2: Fitting curve for $a_{0}$, $a_{1}$, and $a_{2}$ of the 58M model on ID-Val. $R^{2}$ for $a_{0}$ and $a_{1}$ is affected by fluctuations.
From the Dynamic Hyperbolic-law to the Overall Temporal Pattern. In the dynamic hyperbolic-law, $a_{2}$ is the converging factor denoting the converged value of loss on each tail token as context length increases. $a_{0}$ denotes the loss gap between the first token and the last token, and $a_{1}$ is the scaling factor along sequence length. To analyze how the derived values of $a_{0}$, $a_{1}$, and $a_{2}$ vary as training steps increase, we collect the derived value tuples of $(a_{0},a_{1},a_{2})$ for all evaluated small model checkpoints during pretraining: $[(a_{0}^{N_{1}},a_{1}^{N_{1}},a_{2}^{N_{1}}),\cdots,(a_{0}^{N_{tot}},a_{1}^{N_{tot}},a_{2}^{N_{tot}})]$, where $N_{i}$ is the number of trained tokens at the $i$-th checkpoint, and $N_{tot}$ is the total number of tokens used for pretraining. Then we plot the collected $(a_{0},a_{1},a_{2})$ values on ID-Val w.r.t. the number of tokens trained (denoted as $N$) for the 58M model in Fig. 2. It can be seen that the evolution of $a_{0}$, $a_{1}$, and $a_{2}$ presents an evident pattern that can be fit by common functions. Therefore, it is possible for us to firstly fit the temporal evolution of $a_{0}$, $a_{1}$, and $a_{2}$, and then fill them into Eq. 5 to depict the temporal evolution of the overall test loss.
First, for $a_{0}$ and $a_{1}$, we use a series of common functions to fit their temporal evolution, and eventually choose the following functions that yield the lowest fitting errors:
@@
where the $\{\alpha_{i}\}$ and $\{\beta_{i}\}$ fitting parameters are solved by non-linear least squares. Observed from the original data in Fig. 2, we find that the value of $a_{0}$ and $a_{1}$ generally converges after a period of training, but suffer from fluctuations after converging due to the uncertainty of the first prediction position in a sequence, which has no previous context to make prediction. To mitigate the fluctuations, we define a separation point ${N}_{sep}$:
@@
And we manually stabilize $a_{0}$ and $a_{1}$ after ${N}_{sep}$: